OpenStudy (anonymous):

How many photons are produced in a laser pulse of 0.227 J at 481 nm?

OpenStudy (john_es):

Followgin Planck's formula for the energy of n photons, \[E=nh\frac{c}{\lambda}\] where \[c\approx3\cdot10^8\ m/s\\ h=6.62\cdot10^{-34}\ J\cdot s \\ \lambda =481\cdot 10^{-9}\ m\\ E=0.227\ J\] Then \[n=\frac{\lambda E}{hc}\]Substitute and you'll obtaind the answer.